# Navigating Cookie Consent Violations Across the Globe

### v1

@@ 4.1 Overview
Several tools have automated the interactions with consent notices [ghostery_ghosteryautoconsent_2022, aarhus_university_cavi-auconsent-o-matic_2021, ninja_cookie_ninja_2022]. Khandelwal et al. [khandelwal_cookieenforcer_2022] click every element and then detect cookie banner menus after each click. Bollinger et al. [bollinger2022automating] accept/reject cookie consent by using the GDPR-specific Consent-O-Matic tool [aarhus_university_cavi-auconsent-o-matic_2021]. However, these approaches are less thorough and scalable than what is needed for our evaluation. We wish to audit cookie behavior after setting consent for individual categories, as well as auditing cookie banners that do not pop up.
To remedy this deficiency, ConsentChk activates cookie banner menus by using a cookie-button extractor (Section 4.3) and a menu activator (Section 4.4). When accessing a web page, if a cookie banner menu is not detected, the extractor analyzes HTML elements to extract the candidate cookie banner buttons that may activate a menu. Cookie consent preferences are extracted in two steps: (1) cookie-category consent and (2) cookies in each category. Finally, cookie consent violations are detected and classified (Section 4.5).
@@ 4.2 Implementation Details
To extract the set of cookies during a website visit, ConsentChk extracts the cookies sent to servers via the network debugging functionality of Chrome DevTools Protocol that reports all HTTP(S) requests with associated cookies [google_inc_chrome_2021]. Dynamic analysis is advantageous because it reveals “real” occurrences of cookies rather than finding only potential ones, thus reducing false positives. For example, a website may block the use of 3rd-party cookies by preventing the loading of 3rd-party scripts and frames without removing the cookies. This additional analysis overcomes the limitations of prior works that extract all cookies in the browser regardless of whether cookies were actually transferred to servers [sanchez-rola_can_2019].
We posit that a website should enforce a cookie consent preference and remove/prevent cookie placement if (1) the consent choice is recorded in the browser, (2) the cookie matches, and (3) the web page being visited is within the scope of consent. These conditions ensure the consent preference of a cookie applies only when the user browses the web pages to which the user gave their consent preference. For example, the user’s consent preference given to website a.com does not apply to website b.com, although it applies to subdomains like subpage.a.com. ConsentChk maps the cookie declarations in the cookie banners to the cookies used by the website by matching cookie names and domains. For patterns such as _gaxxx or _ga###, we assume an ‘x’ or ‘#’ to match any single character. However, a single ‘#’ at the end of the declaration matches any alpha-numeric string. For example, the declared cookie name "_gatxxx" matches cookie "_gat123". The scope of a consent preference is defined by the domain of the consent cookie (such as OptanonConsent of OneTrust) or other local storage object which records the user’s consent. Similarly, the domain names of a cookie and a declaration match each other if the declared domain is a suffix of the cookie domain. This domain matching scheme is similar to the standard cookie domain matching specification [barth_http_2011].
@@ 4.3 Preference Button Extractor
To increase the coverage of ConsentChk beyond websites using consent libraries, we create a detector for extracting cookie banner buttons. We define a cookie banner button to be an HTML element that, upon a user click, displays a menu for the user to set the cookie consent. For each web page, the cookie-button extractor finds visible HTML elements that represent a button or link in all iframes contained in the page. We consider a, button, div, and span elements which are commonly used to represent links and buttons [chris_coyier_make_2021, will_bontrager_software_llc_linking_2021]. For div elements, we only select leaf elements to avoid unrelated elements. ConsentChk uses a random forest model with feature groups such as $n$-grams and keywords from HTML attributes such as aria-label, class, id, and inner text. To increase its coverage, the activator attempts to click the top k candidate buttons. Specifically, if clicking a non-cookie button navigates to another page or activates no consent preference menus, ConsentChk returns to the initial URL and tries other buttons. Additional details regarding the datasets, annotation process, classification features, performance metrics, and models used for our cookie button extractor can be found in Appendix A.
@@ 4.5 Cookie Consent Violation Classifier
Each cookie is then monitored and classified using the logical rules set in Table 3. Cookies that are placed even after the tool rejects cookies are classified as Ignored Cookie Rejection Violations. Cookies placed without being declared in the cookie library set by the CMP are Undeclared Cookie Violations. These violations are classified based on the actions performed by the cookie consent exerciser and the cookies declared in the CMPs’ cookie libraries.
@@ 5 Measurement Methodology
Table 4: The most popular CMPs with more than 1% market share on the top 1M websites as reported by BuiltWith [builtwith_privacy_2022].
@@
We report results from the five categories defined by cookie libraries we selected — OneTrust, CookiePro, and Cookiebot— which are some of the most widely used on the Web. Each of OneTrust and Cookiebot has a $>$1% market share while CookiePro has 0.23%. These CMPs were chosen due to their inclusion of both cookie lists and cookie declarations. Table 4 shows the market shares and the satisfied criteria of the CMPs. Since CookiePro is part of OneTrust, we combine it into the reported results of OneTrust.
@@
[…]
@@ 5.1 Measurement Procedure
Next, ConsentChk visits other sub-pages that have hyperlinks on the homepage with URLs matching the domain of the consent cookie to generate cookie traffic and check for cookie consent violations. These sub-pages are chosen randomly to ensure a representative sample of the website’s content is analyzed. This step is important for capturing additional cookies used on subsites [hils_measuring_2020]. ConsentChk rejects all rejectable cookie categories.
@@
The crawler uses a 60-second timeout to load the pages. We found this timeout sufficient to completely load most of the web pages with the fast network of our servers and cloud providers. The crawler uses the Playwright browser automation tool [microsoft_microsoftplaywright-python_2020] to control the Google Chrome web browser and utilizes techniques provided by an automatic browsing plugin to avoid getting detected by bot detection (i.e., Puppeteer and Playwright’s stealth mode) [berstend_puppeteer-extra-plugin-stealth_2021]. These added Playwright functionalities mimic human behavior, bot flagging via headless detection and user agents.
[…]
Of the top 20k websites, 10.8k websites were in English and properly loaded. From this subset, ConsentChk found 1,793 websites with cookie banners and consent settings from our list of CMPs. ConsentChk analyzed the cookie consent behavior of 1,312 sites containing cookie preference menus (see the breakdown in Figure 3). ConsentChk collected information on cookie banner parameters, cookies declared in the CMP, cookies placed on the browser, website traffic and navigated subpages.
@@ 5.2 Personal Information Analysis
Most privacy laws only pertain to personal information – that is data that could be used to either identify an individual or data pertaining to an identifiable individual. As we are primarily interested in studying cookie consent violations, we need a method for differentiating cookies containing personal information or identifiers from those not. For example, some cookies may simply contain a number to indicate a user’s preference for light/dark mode on a site while other cookies may contain the latitude and longitude of a user or a unique tracking identifier. The cookie declaration may also include its purposes that describe the nature of data usage. For example, the cookie loc with domain addthis.com collects the location of users to help addthis.com track their location when its share buttons are clicked. Because of this and the fact that a vast majority of the cookie consent violation types are Undeclared Cookies, we need a system to detect and categorize whether cookies contain personal information, using the cookie’s name, value, and declared purpose. We built a lightweight detector that searches for keywords (e.g., city, postal code, state, country name, and uid) and regex patterns (e.g., IP address, GPS location, and common tracker formats), as well as decoding Base64-encoded cookies. The detector also leverages the zxcvbn [wheeler2016zxcvbn] library for measuring the entropy of the cookie values (to discover tracking UIDs). For those cookie purposes not declared by website developers in the cookie libraries, our detector referenced Cookiepedia [one_trust_cookiepedia_2020] as well as other open-source cookie databases. We opted to use this lightweight approach for detecting personal information to scale our approach to millions of cookies.
@@ 5.3 Analysis Methodology
We analyzed 3 different aspects of cookie banners on the websites across the 8 regions: (1) the cookies placed on each website across each region; (2) the differences in cookie consent violations we detected for each cookie on the websites across each region; and (3) the differences in UI settings and parameters for the websites across each region. We report both high-level trends in cookie consent violations and cookie count as well as per-site and pairwise region differences. In reporting total violation counts (Table 6), we report the union across the 10 measurements. In other plots (Figures 5 and 6), we report the average. Unless stated otherwise, the reported results use Ireland as a baseline due to the strict requirements on cookies from the GDPR.
@@ 5.4 Statistical Significance Tests
Table 5: Statistical Significance Tests for Figures 4, 5 and 6.
@@ 6 Measurement Results
[…]
@@ 6.1 High-Level Observations
The results of our crawls demonstrate a stark discrepancy in cookie count, cookie consent violations, and cookie banner implementations across different regions. These differences result in up to 3.44% more websites with undeclared cookies in some regions than the baseline EU and UK regions. The US contains 250,447, almost 3×3$\times$3 × more cookies than websites in the EU, and as much as 10% more undeclared cookies per website. A large number of websites also contain differences in cookie library implementations, with non-European regions containing more cookie banners that do not display “Reject All Cookies” buttons, have shorter consent lifetimes, and have opt-out or implied consent rather than opt-in banners.
@@
We analyze the contents of each cookie found in a cookie consent violation to determine whether they contain personal information, potentially violating user consent. The 75% of cookie consent violations likely involve cookies containing personal information. Table 7 contains statistics on cookies with tracking IDs (64.10%), location (7.44%), IP address (5.74%), and language (1%).
@@ 6.2 Major Findings
Finding 3: 3rd-party cookies still constitute a large number of cookie consent violations across regions. While the EU and UK have significantly fewer 3rd-party cookies (Figure 4(c)), all other regions have a large number of 3rd-party cookies placed on each site. Third-party cookies in the US made up roughly 29.03 cookies per site while websites in the EU and UK contained the fewest with an average of 8.88 and 9.16 cookies per site, respectively. These violations were found to be statistically significant (Section 5.4).
Finding 4: Regions with fewer undeclared cookie consent violations have more ignored cookie rejection and wrong cookie category violations. Figure 5(a) and fig. 5(b) show an inverse correlation between the number of ignored rejection violations and undeclared cookie consent violations. Most likely, website developers are more vigilant about categorizing undeclared cookies for GDPR compliance but neglect to update their non-GDPR CMP cookie lists. Consequently, sites in the EU and UK have higher rates of ignored cookie rejections likely due to an increase in declared cookies, but a lack of properly implemented CMP scripts. Undeclared cookies are placed regardless of rejection. These cookies are not even present in the cookie library and thus cannot be rejected.
Finding 5: Cookie placement and violation rates can be categorized into three groups: EU/UK (fewest cookies, highest compliance), CA, AU, SG, CAN, ZA (moderate cookie count and compliance), US (most cookies, lowest compliance). Without any privacy regulations, the US has poor cookie practices. In our analysis of websites’ cookies and cookie consent violations, we ran pairwise comparisons on websites between different regions, discovering that all other regions, particularly the US, have a significant increase in cookie count and cookie consent violations compared to the EU and UK (Figures 8 and 9). These discrepancies were found to be statistically significant (Section 5.4).
@@ 6.3 Contextualizing Findings with Prior Work
Early audits reported very high violation prevalence even on small samples. Sanchez-Rola (2019) et al. [sanchez-rola_can_2019] found that 92% of 2k audited sites set tracking cookies before consent. Bollinger (2022) et al. [bollinger2022automating] observed an increase of violations to 94.7% of 30k sites. Our crawl found at least one violation on 96.18–97.72% of sites — an absolute increase of 1–3% every three years. Though this increase may be due to many factors: website developers not updating CMPs and repeated measurements catching more violation instances, cookie consent violations appear to be increasing.
@@
Bollinger et al. [bollinger2022automating] reported 82.5% of sites deploying undeclared cookies in the EU. For other non-EU regions, we found that undeclared cookie violations were even more prevalent. Undeclared cookies remained the dominant category, with 90.19–94.15% of sites containing at least one undeclared cookie. These undeclared cookie violations are found to be more prevalent (3–4%) in non-EU regions than EU regions.
@@
Implicit consent —– often manifested as cookies being set despite explicit rejection —– has persisted and even grown more widespread. Bollinger (2022) et al. [bollinger2022automating] report that 69.7% of sites commit implicit-consent violations and 21.3% ignore users’ reject clicks, while Sanchez-Rola (2019) et al. [sanchez-rola_can_2019] found 92% of 2,000 audited sites performed tracking before any consent. Bouhoula (2023) et al. [bouhoula2023automated] show 77.5% of EU sites implicitly consent when banners are closed. We found that “Ignored Cookie Rejections” affect 80.74% (Australia) up to 88.41% (California) of sites — with 44,605 ignored-rejection cookies in the UK and 128,871 in the US.
@@
Rasaii et al. [rasaii2023exploring] reported banner prevalence of 47% in EU vs. 30% non-EU, and 83-–96% more tracking cookies outside the EU. Eijk et al. [eijk_impact_2019] measured a 102% increase in banner odds in the EU. We extend these findings by quantifying both violation counts and UI configurations: US sites average 46.5 first-party and 37.3 third-party cookies per site, approximately three times the EU’s 20.0 and 11.0. Meanwhile, non-EU regions are 3.4% more likely to omit “Reject All” buttons and default to opt-out or implied consent models (32% of US sites).

### v2

@@ 4.1 Overview
Several tools have automated the interactions with consent notices [31, 68, 47]. Khandelwal et al. [39] click every element and then detect cookie banner menus after each click. Bollinger et al. [9] accept/reject cookie consent by using the GDPR-specific Consent-O-Matic tool [68]. However, these approaches are less thorough and scalable than what is needed for our evaluation. We wish to audit cookie behavior after setting consent for individual categories, as well as auditing cookie banners that do not pop up.
To remedy this deficiency, ConsentChk activates cookie banner menus by using a cookie-button extractor (Section˜4.3) and a menu activator (Section˜4.4). When accessing a web page, if a cookie banner menu is not detected, the extractor analyzes HTML elements to extract the candidate cookie banner buttons that may activate a menu. Cookie consent preferences are extracted in two steps: (1) cookie-category consent and (2) cookies in each category. Finally, cookie consent violations are detected and classified (Section˜4.5).
@@ 4.2 Implementation Details
To extract the set of cookies during a website visit, ConsentChk extracts the cookies sent to servers via the network debugging functionality of Chrome DevTools Protocol that reports all HTTP(S) requests with associated cookies [36]. Dynamic analysis is advantageous because it reveals “real” occurrences of cookies rather than finding only potential ones, thus reducing false positives. For example, a website may block the use of 3rd-party cookies by preventing the loading of 3rd-party scripts and frames without removing the cookies. This additional analysis overcomes the limitations of prior works that extract all cookies in the browser regardless of whether cookies were actually transferred to servers [56].
We posit that a website should enforce a cookie consent preference and remove/prevent cookie placement if (1) the consent choice is recorded in the browser, (2) the cookie matches, and (3) the web page being visited is within the scope of consent. These conditions ensure the consent preference of a cookie applies only when the user browses the web pages to which the user gave their consent preference. For example, the user’s consent preference given to website a.com does not apply to website b.com, although it applies to subdomains like subpage.a.com. ConsentChk maps the cookie declarations in the cookie banners to the cookies used by the website by matching cookie names and domains. For patterns such as _gaxxx or _ga###, we assume an ‘x’ or ‘#’ to match any single character. However, a single ‘#’ at the end of the declaration matches any alpha-numeric string. For example, the declared cookie name "_gatxxx" matches cookie "_gat123". The scope of a consent preference is defined by the domain of the consent cookie (such as OptanonConsent of OneTrust) or other local storage object which records the user’s consent. Similarly, the domain names of a cookie and a declaration match each other if the declared domain is a suffix of the cookie domain. This domain matching scheme is similar to the standard cookie domain matching specification [6].
@@ 4.3 Preference Button Extractor
To increase the coverage of ConsentChk beyond websites using consent libraries, we create a detector for extracting cookie banner buttons. We define a cookie banner button to be an HTML element that, upon a user click, displays a menu for the user to set the cookie consent. For each web page, the cookie-button extractor finds visible HTML elements that represent a button or link in all iframes contained in the page. We consider a, button, div, and span elements which are commonly used to represent links and buttons [15, 44]. For div elements, we only select leaf elements to avoid unrelated elements. ConsentChk uses a random forest model with feature groups such as $n$-grams and keywords from HTML attributes such as aria-label, class, id, and inner text. To increase its coverage, the activator attempts to click the top k candidate buttons. Specifically, if clicking a non-cookie button navigates to another page or activates no consent preference menus, ConsentChk returns to the initial URL and tries other buttons. Additional details regarding the datasets, annotation process, classification features, performance metrics, and models used for our cookie button extractor can be found in Appendix A.
@@ 4.5 Cookie Consent Violation Classifier
Each cookie is then monitored and classified using the logical rules set in Table˜3. Cookies that are placed even after the tool rejects cookies are classified as Ignored Cookie Rejection Violations. Cookies placed without being declared in the cookie library set by the CMP are Undeclared Cookie Violations. These violations are classified based on the actions performed by the cookie consent exerciser and the cookies declared in the CMPs’ cookie libraries.
@@ 5 Measurement Methodology
Table 4: The most popular CMPs with more than 1% market share on the top 1M websites as reported by BuiltWith [13].
@@
We report results from the five categories defined by cookie libraries we selected — OneTrust, CookiePro, and Cookiebot— which are some of the most widely used on the Web. Each of OneTrust and Cookiebot has a >>1% market share while CookiePro has 0.23%. These CMPs were chosen due to their inclusion of both cookie lists and cookie declarations. Table 4 shows the market shares and the satisfied criteria of the CMPs. Since CookiePro is part of OneTrust, we combine it into the reported results of OneTrust.
@@
From the top 20k global websites in the Tranco list November 2023 (ID: 5Y3LN), we select $\numprint{10436}$ websites, which have an English homepage and were loaded successfully, for further analysis with ConsentChk. Some websites in the list failed to load for various reasons. For example, some URLs are non-website ad-serving domains. The language of the websites is determined by a neural-network-based language detector after converting the web pages to plain text [32, 59].
@@ 5.1 Measurement Procedure
Next, ConsentChk visits other sub-pages that have hyperlinks on the homepage with URLs matching the domain of the consent cookie to generate cookie traffic and check for cookie consent violations. These sub-pages are chosen randomly to ensure a representative sample of the website’s content is analyzed. This step is important for capturing additional cookies used on subsites [35]. ConsentChk rejects all rejectable cookie categories.
@@
The crawler uses a 60-second timeout to load the pages. We found this timeout sufficient to completely load most of the web pages with the fast network of our servers and cloud providers. The crawler uses the Playwright browser automation tool [46] to control the Google Chrome web browser and utilizes techniques provided by an automatic browsing plugin to avoid getting detected by bot detection (i.e., Puppeteer and Playwright’s stealth mode) [7]. These added Playwright functionalities mimic human behavior, bot flagging via headless detection and user agents.
We conducted experiments in a distributed framework based on Docker Swarm [37] on 4 machines with 1.08TB RAM and 96 task queue workers. The cookie-consent scanning of the $\numprint{10436}$ websites took 40 hours to perform the measurements from the 8 locations. The crawls were performed during October 4–12, 2024.
Of the top 20k websites, 10.8k websites were in English and properly loaded. From this subset, ConsentChk found 1,793 websites with cookie banners and consent settings from our list of CMPs. ConsentChk analyzed the cookie consent behavior of 1,312 sites containing cookie preference menus (see the breakdown in Figure˜3). ConsentChk collected information on cookie banner parameters, cookies declared in the CMP, cookies placed on the browser, website traffic and navigated subpages.
@@ 5.2 Personal Information Analysis
Most privacy laws only pertain to personal information – that is data that could be used to either identify an individual or data pertaining to an identifiable individual. As we are primarily interested in studying cookie consent violations, we need a method for differentiating cookies containing personal information or identifiers from those not. For example, some cookies may simply contain a number to indicate a user’s preference for light/dark mode on a site while other cookies may contain the latitude and longitude of a user or a unique tracking identifier. The cookie declaration may also include its purposes that describe the nature of data usage. For example, the cookie loc with domain addthis.com collects the location of users to help addthis.com track their location when its share buttons are clicked. Because of this and the fact that a vast majority of the cookie consent violation types are Undeclared Cookies, we need a system to detect and categorize whether cookies contain personal information, using the cookie’s name, value, and declared purpose. We built a lightweight detector that searches for keywords (e.g., city, postal code, state, country name, and uid) and regex patterns (e.g., IP address, GPS location, and common tracker formats), as well as decoding Base64-encoded cookies. The detector also leverages the zxcvbn [69] library for measuring the entropy of the cookie values (to discover tracking UIDs). For those cookie purposes not declared by website developers in the cookie libraries, our detector referenced Cookiepedia [67] as well as other open-source cookie databases. We opted to use this lightweight approach for detecting personal information to scale our approach to millions of cookies.
@@ 5.3 Analysis Methodology
We analyzed 3 different aspects of cookie banners on the websites across the 8 regions: (1) the cookies placed on each website across each region; (2) the differences in cookie consent violations we detected for each cookie on the websites across each region; and (3) the differences in UI settings and parameters for the websites across each region. We report both high-level trends in cookie consent violations and cookie count as well as per-site and pairwise region differences. In reporting total violation counts (Table˜6), we report the union across the 10 measurements. In other plots (Figures˜5 and 6), we report the average. Unless stated otherwise, the reported results use Ireland as a baseline due to the strict requirements on cookies from the GDPR.
@@ 5.4 Statistical Significance Tests
Table 5: Statistical Significance Tests for Figures˜4, 5 and 6.
@@ 6 Measurement Results
Websites made it challenging for users to opt out of unnecessary cookies. Despite the Necessary cookie category being designed to indicate cookies required purely for website services, we found 9.81% of websites with un-Necessary cookie categories set unnecessary cookies as "always active." This would prevent users from opting out of tracking practices.
@@ 6.1 High-Level Observations
The results of our crawls demonstrate a stark discrepancy in cookie count, cookie consent violations, and cookie banner implementations across different regions. These differences result in up to 3.44% more websites with undeclared cookies in some regions than the baseline EU and UK regions. The US contains 250,447, almost $3\times$ more cookies than websites in the EU, and as much as 10% more undeclared cookies per website. A large number of websites also contain differences in cookie library implementations, with non-European regions containing more cookie banners that do not display “Reject All Cookies” buttons, have shorter consent lifetimes, and have opt-out or implied consent rather than opt-in banners.
@@
We analyze the contents of each cookie found in a cookie consent violation to determine whether they contain personal information, potentially violating user consent. The 75% of cookie consent violations likely involve cookies containing personal information. Table˜7 contains statistics on cookies with tracking IDs (64.10%), location (7.44%), IP address (5.74%), and language (1%).
@@ 6.2 Major Findings
Finding 3: 3rd-party cookies still constitute a large number of cookie consent violations across regions. While the EU and UK have significantly fewer 3rd-party cookies (Figure˜4(c)), all other regions have a large number of 3rd-party cookies placed on each site. Third-party cookies in the US made up roughly 29.03 cookies per site while websites in the EU and UK contained the fewest with an average of 8.88 and 9.16 cookies per site, respectively. These violations were found to be statistically significant (Section˜5.4).
Finding 4: Regions with fewer undeclared cookie consent violations have more ignored cookie rejection and wrong cookie category violations. Figure˜5(a) and fig.˜5(b) show an inverse correlation between the number of ignored rejection violations and undeclared cookie consent violations. Most likely, website developers are more vigilant about categorizing undeclared cookies for GDPR compliance but neglect to update their non-GDPR CMP cookie lists. Consequently, sites in the EU and UK have higher rates of ignored cookie rejections likely due to an increase in declared cookies, but a lack of properly implemented CMP scripts. Undeclared cookies are placed regardless of rejection. These cookies are not even present in the cookie library and thus cannot be rejected.
Finding 5: Cookie placement and violation rates can be categorized into three groups: EU/UK (fewest cookies, highest compliance), CA, AU, SG, CAN, ZA (moderate cookie count and compliance), US (most cookies, lowest compliance). Without any privacy regulations, the US has poor cookie practices. In our analysis of websites’ cookies and cookie consent violations, we ran pairwise comparisons on websites between different regions, discovering that all other regions, particularly the US, have a significant increase in cookie count and cookie consent violations compared to the EU and UK (Figures˜8 and 9). These discrepancies were found to be statistically significant (Section˜5.4).
@@ 6.3 Contextualizing Findings with Prior Work
Early audits reported very high violation prevalence even on small samples. Sanchez-Rola (2019) et al. [56] found that 92% of 2k audited sites set tracking cookies before consent. Bollinger (2022) et al. [9] observed an increase of violations to 94.7% of 30k sites. Our crawl found at least one violation on 96.18–97.72% of sites — an absolute increase of 1–3% every three years. Though this increase may be due to many factors: website developers not updating CMPs and repeated measurements catching more violation instances, cookie consent violations appear to be increasing.
@@
Bollinger et al. [9] reported 82.5% of sites deploying undeclared cookies in the EU. For other non-EU regions, we found that undeclared cookie violations were even more prevalent. Undeclared cookies remained the dominant category, with 90.19–94.15% of sites containing at least one undeclared cookie. These undeclared cookie violations are found to be more prevalent (3–4%) in non-EU regions than EU regions.
@@
Implicit consent —– often manifested as cookies being set despite explicit rejection —– has persisted and even grown more widespread. Bollinger (2022) et al. [9] report that 69.7% of sites commit implicit-consent violations and 21.3% ignore users’ reject clicks, while Sanchez-Rola (2019) et al. [56] found 92% of 2,000 audited sites performed tracking before any consent. Bouhoula (2023) et al. [10] show 77.5% of EU sites implicitly consent when banners are closed. We found that “Ignored Cookie Rejections” affect 80.74% (Australia) up to 88.41% (California) of sites — with 44,605 ignored-rejection cookies in the UK and 128,871 in the US.
@@
Rasaii et al. [54] reported banner prevalence of 47% in EU vs. 30% non-EU, and 83-–96% more tracking cookies outside the EU. Eijk et al. [28] measured a 102% increase in banner odds in the EU. We extend these findings by quantifying both violation counts and UI configurations: US sites average 46.5 first-party and 37.3 third-party cookies per site, approximately three times the EU’s 20.0 and 11.0. Meanwhile, non-EU regions are 3.4% more likely to omit “Reject All” buttons and default to opt-out or implied consent models (32% of US sites).